An Apple annual coupon bond has a coupon rate of 5.1% face value of $1000, and 4 years to maturity. If its yield to maturity is 5.1%, what is its Macaulay duration? answer in years, rounded to three decimal places
Answer #1

Macaulay Duration = 3.718 Years

Explanation

Macaulay Bond duration is nothing but the weighted average of time where weights are taken from present value of cash flow generated by the bonds.

    t CF DF PV W    t * W
1 51 0.951475 48.52521 0.048525 0.048525
2 51 0.905304 46.17052 0.046171 0.092341
3 51 0.861374 43.93008 0.04393 0.13179
4 1051 0.819576 861.3742 0.861374 3.445497
1000 3.718153

t = Years on time line.

CF = Cash flow. As we know during the 3 years only cash flow is the coupon amount = 1000 * 5.10 % = 51

But in the year - 4 ........ we get the face value paid back = 1000 + 51 = 1051

Discounting factors are computed as ....... 1/1.051 = 0.951475 and there after continue to press " = " to get remaining years DF values.

PV = CF * DF

W = weights are obtained as PV / 1000

t * W is the product of first column ( t) with ( W)

Total of ( t * W ) .........gives the bond duration.